How do you find the extent of the graph of a hyperbola in both forms? (x^2/a^2 - y^2/b^2 = 1, and y^2/a^2 - x^2/b^2 = 1)
Answer by lwsshak3(11628)

How do you find the extent of the graph of a hyperbola in both forms? (x^2/a^2 - y^2/b^2 = 1, and y^2/a^2 - x^2/b^2 = 1)
..
What you have listed are the 2 standard forms of an hyperbola.
The first form, when the x^2 term is listed first, the transverse axis is horizontal, that is, it opens sideways. Conversely, when the y^2 term is listed first, the transverse axis opens upwards. Other important information which you will need to graph a hyperbola can be found in any algebra textbook. They are listed below for the second type (y^2 term first) of hyperbolas with centers at (0,0).
Foci(0+-c),c^2=a^2+b^2
Vertices (0+-a)
Asymptotes y==-(a/b)x
Length of transverse axis (2a)
Length of conjugate axis (2b)
Eccentricity e=c/a
a>b,a=b,or a less than b